OpenStudy (anonymous):

Find the indefinite integral: Integrand of 16x((8x^2)+7)^6

sam (.sam.):

U-sub, u=8x^2+7 du=16x dx \[\int\limits u^6 \, du\] Then integrate

OpenStudy (anonymous):

so my answer would be 1/9(8x^2+7)^9? The last exponent was supposed to be an 8, not a 6 (sorry)

OpenStudy (anonymous):

+ C at the end of course

sam (.sam.):

Check your integration, it should be \[\frac{u^7}{7}+c\]

OpenStudy (anonymous):

got it thanks!

sam (.sam.):

welcome :)